OpenStudy (anonymous):

Use logarithmic differentiation to find the derivative of the function y=x^(1/x)

OpenStudy (anonymous):

not sure how to take derivative of (1/x)lnx

OpenStudy (anonymous):

\[ \ln(x^\frac{1}{x}) = \frac{\ln(x)}{x} \]

OpenStudy (anonymous):

You can use quotient rule.

OpenStudy (cwrw238):

use the product rule for (1/x) ln x

OpenStudy (anonymous):

-1x^-2(lnx)+x^-1(1/x)

OpenStudy (cwrw238):

yes - quotient rule will do it also

OpenStudy (anonymous):

\[ \frac{d}{dx}\frac{f}{g} =\frac{ f'g-fg'}{g^2}\]

OpenStudy (anonymous):

did i do the product rule right?

OpenStudy (anonymous):

oh. got it. thanks

OpenStudy (cwrw238):

yes